Alex21111:

You can stuff envelopes twice as fast as your friend. Together you can stuff 6750 envelopes in 4.5 hours. How long will it take each of you working alone?

planetpeebs:

i think u have to add. im not completl sure @extrinix

planetpeebs:

@snowflake0531 u to

planetpeebs:

the dude who wanted this answered is offline --.--

SmokeyBrown:

Let's say the "number of envelopes I can stuff in an hour" is a, and the "number of envelopes my friend can stuff in an hour" is b. Since I stuff envelopes twice as fast, we can say that a = 2b Also, the number of envelopes we can both stuff (in 4.5 hours) is 6750. We'll represent this with the equation 4.5(a+b) = 6750. Use these two equations to find out how many envelopes I can stuff in one hour (a) and how many my friend can (b). Then, it's simple to find out how many hours it takes each of you working alone to stuff 6750

Florisalreadytaken:

@smokeybrown wrote:
Let's say the "number of envelopes I can stuff in an hour" is a, and the "number of envelopes my friend can stuff in an hour" is b. Since I stuff envelopes twice as fast, we can say that a = 2b
it would be \[ a=\frac{1}{2}b \] or \[ 2a=b \]

SmokeyBrown:

I think it is a = 2b, since the number of envelopes 'a' should be twice as many as the number of envelopes 'b'. For example, 'a' could be 200 and 'b' could be 100, since the number is twice as much; then the equation would make sense: 200 = 2*100

Florisalreadytaken:

yes, fair enough